OpenStudy (sophadof):

Equation Below Fan and a Medal!!(:

OpenStudy (sophadof):

\[-\frac{ z-2 }{ 10z-15 }=\frac{ 1 }{ 2z-3 }-\frac{ 6 }{ 2x-3 }\]

OpenStudy (sophadof):

@mathmale

OpenStudy (mathmale):

Hello! What kind of help do you need? I see that one of your denominators ix 2z-3 and another is 2x-3. Is that correct?

OpenStudy (sophadof):

Sorry @mathmale they are all x's

OpenStudy (sophadof):

that was a typo

OpenStudy (mathmale):

Please check this over and see if it exactly reproduces the problem printed in your materials. \[-\frac{ x-2 }{ 10x-15 }=\frac{ 1 }{ 2x-3 }-\frac{ 6 }{ 2x-3 }\]

OpenStudy (sophadof):

dang.. it I meant they are all z's, I am having a tough time right now focusing well.. But yes that is correct but with all z's @mathmale

OpenStudy (mathmale):

Let's stay with all x's so long as the equation is correct otherwise.

OpenStudy (sophadof):

All right!!

OpenStudy (mathmale):

Please express in y our own words what the instructions for this problem say.

OpenStudy (sophadof):

I thought this would be not possible because they all have the same letter variable.. @mathmale

OpenStudy (mathmale):

Let's go with the flow; I think you'll see that there are now no obstacles. But again, what is your objective here? What are you supposed to do?

OpenStudy (sophadof):

I am supposed to solve for z(in this case x) @mathmale

OpenStudy (mathmale):

Just change that to, "I am supposed to solve for x." ;) take a good, hard look at the 3 denominators. Do they seem to have anything in common? If so, explain what they have in common.

OpenStudy (sophadof):

I am supposed to solve for x. They can all be multiplied by 2 numbers? @mathmale

OpenStudy (mathmale):

Could you please be a bit more specific?: Why multiply some or all of the denominators "by 2 numbers"? First, name your objective. Why would we want to multiply any of the den. by something else? Note: you are not wrong...I just want more specifics.

OpenStudy (mathmale):

What's a general rule for combining fractions, by the way?

OpenStudy (sophadof):

I really don't know :( I am very bad at math and I am really hard on myself about it... I am trying to learn but I just don't know @mathmale

OpenStudy (mathmale):

Look at the denominators of the middle and right fractions. They're the same, correct?

OpenStudy (sophadof):

Correct @mathmale

OpenStudy (mathmale):

Look at the denom. of the leftmost fraction. Could it be factored? If so, what are the factors?

OpenStudy (sophadof):

5x-3? @mathmale

OpenStudy (mathmale):

Hint: Is there any way in which you could factor this denominator so that one of the factors is the same as the factor of the right two fractions?

OpenStudy (sophadof):

sorry! I meant 2x-3! I divided by 5 then had the number 5 stuck in my mind :D @mathmale

OpenStudy (mathmale):

So, your first den. (denominator) is 5(2x-3). The 5 is unique to that den.; the other two dens. don't have it as a factor. To combine these fractions, or to eliminate the dens. altogether, we need to have exactly the same den. in each fraction. How would you accomplish that?

OpenStudy (sophadof):

would you combine the numbers that have the x's? @mathmale

OpenStudy (mathmale):

Let's copy down the equation as it now stands: }\[-\frac{ x-2 }{ 10x-15 }=\frac{ 1 }{ 2x-3 }-\frac{ 6 }{ 2x-3 }\]

OpenStudy (mathmale):

As you said, we can factor 5 out of the denom. 10x-15, resulting in our having 5(2x-3). We could have all the dens. the same if we'd multiply both numerator and denominator of the 2 rightmost fractions by 5. Would y ou please try that?

OpenStudy (mathmale):

If you've done this correctly, all of your fractions will have the same den.

OpenStudy (mathmale):

I will be done for about 2 minutes. Continue working on this problem, please.

OpenStudy (sophadof):

I will continue! @mathmale

OpenStudy (mathmale):

don't spend a lot of time on this. All you really have to do is to put parentheses around each numerator in the 2nd 2 fractions and the same around each den., and then multiply both num. and den. by 5. That's it. Don't multiply beyond that, at least not yet.

OpenStudy (mathmale):

How are you doing on this work?

OpenStudy (mathmale):

@sophadof?

OpenStudy (sophadof):

It's confusing me... @mathmale sorry i took so long to respond

OpenStudy (sophadof):

so then all the denominators look like this? 5(2x-3) @mathmale

OpenStudy (mathmale):

Indeed they do. What are the new numerators for the right 2 fractions?

OpenStudy (sophadof):

That would be...

OpenStudy (sophadof):

5 and 30? @mathmale

OpenStudy (mathmale):

After multiplying numerator and den. of both of the right two fractions by 5, we should have:\[-\frac{ x-2 }{ 10x-15 }=\frac{5* 1 }{5*( 2x-3) }-\frac{5* 6 }{5( 2x-3) }\]

OpenStudy (mathmale):

Yes, 5 and 30. Very good.

OpenStudy (mathmale):

Now, since all of the dens. are the same, we can simply cross out the dens. This leaves us with the numerators only, in the following equation: -(x-2)=5-30

OpenStudy (mathmale):

Please simplify this, and solve for x.

OpenStudy (mathmale):

\[-\frac{ x-2 }{ 10x-15 }=\frac{ 5*1 }{5( 2x-3) }-\frac{5( 6) }{5( 2x-3) }\]

OpenStudy (sophadof):

27!? @mathmale

OpenStudy (mathmale):

that looks good! Ideally, you'd check your result, x=27, to ensure that it truly does make the given equation true.

OpenStudy (sophadof):

OH MY GOSH! THANK YOU SO SO SO MUCH! (: @mathmale

OpenStudy (mathmale):

If yes, then x=27 or z=27 is your solution. My great pleasure to work with you! See you again.